OpenStudy (anonymous):

find the product: (x-4)(x+4)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Are you still there @jdabaddest?

OpenStudy (skullpatrol):

Have you seen the special case of the distributive property @jdabaddest?

OpenStudy (skullpatrol):

$$\Huge (a+b)(a-b)=a^2-b^2$$ $$\Huge (x+4)(x-4)=?$$